OpenStudy (anonymous):

A cylindrical vase has a diameter of 4 inches. At the bottom of the vase, there are 6 marbles, each of diameter 2 inches. The vase is filled with water up to a height of 8 inches. What is the volume of water in the vase? 26π in^3 16π in^3 24π in^3 32π in^3

OpenStudy (anonymous):

@jim_thompson5910 @amistre64

jimthompson5910 (jim_thompson5910):

Total Volume = (Volume of Cylinder) - (Volume of Marbles)

jimthompson5910 (jim_thompson5910):

Volume of Cylinder V = pi*r^2*h V = pi*2^2*8 V = 32pi

jimthompson5910 (jim_thompson5910):

Volume of One Marble: V = (4/3)*pi*r^3 V = (4/3)*pi*1^3 V = (4/3)*pi Volume of 6 Marbles V = 6*(Volume of one marble) V = 6*((4/3)*pi) V = 8pi

jimthompson5910 (jim_thompson5910):

So.... Total Volume = (Volume of Cylinder) - (Volume of Marbles) Total Volume = (Volume of Cylinder) - (Volume of 6 Marbles) Total Volume = ... I'm sure you know what to do from here

OpenStudy (anonymous):

yes thanks :)

jimthompson5910 (jim_thompson5910):

np